When temperature of a gas is 20°C and pressure is changed from P1 = 1.01*105 Pa to P2 =1.165 * 105 Pa then the volume changed by 10%.The bulk modulus is
        (a) 1.55 * 105 Pa
        (b) 0.115 * 105 Pa
        (c) 1.4 * 105 Pa
        (d) 1.01 * 105 Pa

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol :- (a) B= ∆p/∆V/V = (1.165 * 105 – 1.01*105)/0.1
= 1.55 * 105 Pa